Please do not give final answer to equation.
Am having trouble finding correct steps to rewriting following equation without fractions.
Equation: 1 - 3/x - 28/x^2 = 0
Is LCD just (x), or (x)(x^2)?
Thanks

Answer by ewatrrr(24785)

 
Hi,
LCD is x^2
x^2 - 3x - 28 = 0
(x+4)(x-7) = 0